OpenStudy (anonymous):

what is the equation written in slope-intercept form that passes through the point (-1,-6) and is perpendicular to the line y+6=-1/4(x-9)

OpenStudy (anonymous):

@Godlovesme

OpenStudy (godlovesme):

first write y+6= -1/4(x-9) in slope-intercept formula to do that distribute -1/2 first and subtract 6 can u do that?

OpenStudy (anonymous):

do you mean distribute -1/4?

OpenStudy (godlovesme):

@priea

OpenStudy (godlovesme):

yes i apologize =0

OpenStudy (anonymous):

its okay just making sure i understand

OpenStudy (godlovesme):

k so can u write the eqn in slope-intercept ? :)

OpenStudy (anonymous):

yea almost but for some reason i got 2.25 when i distributed -1/4 with -9. i think i did something wrong

OpenStudy (godlovesme):

no no u r right its y+6= -1/4x + 2.25 next step subtract 6

OpenStudy (anonymous):

y=-1/4x-3.75?

OpenStudy (godlovesme):

yes do u remember what we do for the slope when its perpendicular?

OpenStudy (anonymous):

flip it?

OpenStudy (godlovesme):

yes and change the sign so our new slope is?

OpenStudy (anonymous):

thats where im lost

OpenStudy (godlovesme):

k the slope was -1/4 flip --> -4 /1 change sign --> 4/1 --> 4 does that make sense?

OpenStudy (anonymous):

so y=4x-3.75?

OpenStudy (godlovesme):

yes sorry for taking so long my internet slow (-1,-6) now we need to use the point-slope form y-y1= m(x-x1) y- -6 = 4(x- -1) y+6= 4(x+1) can u solve from there? :)

OpenStudy (anonymous):

thanks

OpenStudy (godlovesme):

anytime :)

OpenStudy (anonymous):

okay so i really cant fail this lesson anymore so i like need fast help and answers! lol please

OpenStudy (godlovesme):

k what's ur question?

OpenStudy (anonymous):

what is the equation of the line written in slope-intercept form that passes through (3,-3) and is parallel to the line y+2=2(x-3)

OpenStudy (godlovesme):

can u at least change it to slope-intercept form (y=mx+b) all u gotta do is distribute and subtract two

OpenStudy (anonymous):

i thught it was y-y1=m(x-x1)? so then it would be y-2=2(x-3)/

OpenStudy (godlovesme):

yea u r right but first the eqn y+2=2(x-3) should be changed into y=mx+b form that way we will be able to identify the slope sorry for confusing u :(

OpenStudy (anonymous):

sorry laptop died but the answer is y=2x-9?

OpenStudy (anonymous):

OMG thnks so much for your help!

OpenStudy (godlovesme):

yes and my pleasure :D